Grade 11Algebra

HI was heated in the sealed tube at 400 degree Celsius still the equilibrium .HIHwas found to be 22% decomposed the equilibriumCONSTANT of the dissociation is

Sudhanva G V
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
Dear student
 
 

there r 2of eq. constant

Kc  & Kp  

           rxn       2HI -> H2 +I2

  moles initally   2        0     0

moles @eqlb.   2-2x    x      x    

    

                                Kc   =  x*x/{2-2x}^2 = (0.22)^2/4*(0.78)^2 = 0.19888

 

conc.(C)= P/RT  for gas phase rxn  

                                         RT *(x)= PH2 = PI2   = 0.19888        (same as   Kc bcz  power o RT cancel out  )